# Bath-induced interactions and transient dynamics in open quantum systems at strong coupling: Effective Hamiltonian approach

### v1

@@ I Introduction
There exists, in turn, a present interest in understanding environmental effects in quantum systems. Such is the case in, for instance, biological systems and light-harvesting complexes Duan et al. (2022). Emerging technologies such as thermal machines Kosloff and Levy (2014); Goold et al. (2016); Benenti et al. (2017); Mitchison (2019); Josefsson et al. (2018); Ronzani et al. (2018); Mosso et al. (2019), which operate in the quantum regime, also require a deep understanding of system-environment effects for efficient operation. Moreover, light-matter interactions between the vacuum fluctuations and a material placed inside the cavity provides a new scheme in controlling/engineering material/molecular properties Garcia-Vidal et al. (2021); Hübener et al. (2021); Owens et al. (2022); Schlawin et al. (2022); Bloch et al. (2022); Galego et al. (2015); Herrera and Spano (2016). An important remark is that in most of these cases, the system-environment interaction energy cannot be regarded as an asymptotically weak parameter in the configuration. This limits the applicability of standard perturbative open quantum systems theory Breuer and Petruccione (2007).
@@
Starting from out-of-equilibrium initial conditions, it is also of interest to understand the transient dynamics that lead the system and bath to the state of thermal equilibrium. In this respect, many impressive techniques have come to fruition in the last three decades to target the problem. These include; for example, path integral methods Makri (1995); Segal et al. (2010); Kilgour et al. (2019); Kundu and Makri (2023); Strathearn et al. (2018); Gribben et al. (2020), hierarchical equations of motion (HEOM) Tanimura (2020); Lambert et al. (2023), Monte Carlo approaches Erpenbeck et al. (2023); Härtle et al. (2015); Cohen et al. (2015), techniques based on machine learning calculations Luo et al. (2022), tensor networks Prior et al. (2010); Strathearn et al. (2018); Brenes et al. (2020); Cygorek et al. (2022), generalized quantum master equations Shi and Geva (2003); Zhang et al. (2006); Kelly et al. (2016), and diverse Markovian embeddings Woods et al. (2014); Chin et al. (2010); Imamoglu (1994); Garraway (1997a, b); Menczel et al. (2024); all of which can be applied efficiently to study dynamics of system-bath configurations depending on the nature of the microscopic model, energy scales, and computational costs.
@@
The effective Hamiltonian theory Anto-Sztrikacs et al. (2023a, b) has been presented as an alternative to fully numerical techniques. Previous works, focused on the equilibrium and steady-state behavior of impurity systems coupled to bosonic thermal reservoirs showed that this theory can provide an understanding of strong coupling effects in thermal and electronic devices without assumptions on the energy scale of the system-bath interaction energy. The theory relies on a Markovian embedding with a subsequent unitary transformation that imprints environmental effects onto the system’s degrees of freedom. One then considers a specific sub-manifold of the resulting Hamiltonian, and such an approximation can be justified under certain assumptions on the temperature and the bath’s spectral function. As a whole, the effective Hamiltonian mapping prepares a dressed system Hamiltonian that embeds strong coupling effects. The effective system itself is assumed to couple to a residual bath in a Born-Markov limit Breuer and Petruccione (2007). With this method, one can obtain analytical expressions for equilibrium states in the weak, intermediate, and strong system-bath coupling regimes Anto-Sztrikacs et al. (2023b). Furthermore, the technique can be employed even for complex quantum systems that may display phase transitions induced by collective or individual baths Min et al. (2023).
[…]
@@
In Sec. II we describe our models, the theory of effective Hamiltonian, and the Redfield quantum master equation (QME) employed to describe the transient dynamics of the effect of the bath on a quantum system. Sec. III is devoted to displaying our results on equilibrium properties and dynamical signals. In light of our results, we provide a general assessment of the effective Hamiltonian method for both transient or steady-state problems in Sec. IV. Concluding remarks and the experimental relevance of our results are included in Sec. V.
II Dynamics and bath-induced interactions
This Section is devoted to the description of the models, methods and techniques employed in our work. In Sec. II.1, we describe the open quantum system model we consider here. In Sec. II.2, we present the analytical effective Hamiltonian treatment used to study its dynamics. The simulation method we use is explained in Sec. II.3. We also perform a spectral analysis of the effective Hamiltonian in Sec. II.4. This analysis is relevant for understanding the dynamics of out-of-equilibrium configurations.
@@ II.1 Spin-bath models
Figure 1: A two-spin system ($\hat{H}_{S}$, $N=2$) with spin splittings $\Delta_{1}$ and $\Delta_{2}$ is coupled to a thermal bath ($\hat{H}_{B}$) at inverse temperature $\beta$ with a coupling operator $\hat{S}$, modeled by an infinite collection of harmonic modes. The spectral density of the reservoir is given by $J(\omega)$. The spins are not coupled to each other but an effective coupling arises from the interaction with the common bath.
@@
The effect of the environment is typically described via a spectral density function, which may be defined as $J(\omega)=\sum_{k}|t_{k}|^{2}\delta(\omega-\nu_{k})$. The choice of $\hat{S}$ has intricate consequences for the dynamics of the system and, whenever the interaction energy between the system and environment is non-negligible, it will also affect the thermodynamic equilibrium Anto-Sztrikacs et al. (2023a). In this work, we shall focus on the case whereby the coupling to the environment generates coherent evolution towards the equilibration of the system, by choosing $\hat{S}$ along a different component than that of the spin-system Hamiltonian.
@@ II.2 Reaction-coordinate mapping and effective Hamiltonian theory
We want to make minimal assumptions about the system-bath interaction energy and, therefore, we cannot resort to the conventional weak-coupling open systems theory Breuer and Petruccione (2007). For certain spectral density functions of the bath, $J(\omega)$, it is beneficial to employ the reaction-coordinate mapping to go beyond the weak-coupling approximation Martinazzo et al. (2011); Iles-Smith et al. (2014, 2016); Strasberg et al. (2016); Hughes et al. (2009a, b); Anto-Sztrikacs and Segal (2021); Correa et al. (2019); Restrepo et al. (2019).
For the model at hand, we extract a single collective reaction coordinate (RC) from the bath to which the system is strongly coupled. This mapping is useful if the residual bath, i.e., the one composed of the degrees of freedom of the bath remaining after the extraction of the reaction coordinate, is weakly coupled to the enlarged system Hamiltonian. The enlarged system Hamiltonian is composed of the original system and the reaction coordinate. After such mapping, Eq. (1) translates to Anto-Sztrikacs et al. (2023a)
@@
In the above equation, the bosonic operator $\hat{a}$ corresponds to the reaction coordinate with frequency $\Omega$. The parameter $\lambda$ is the coupling between the original system $\hat{H}_{S}$ and the reaction coordinate via the system operator $\hat{S}$. The enlarged system, which is composed of $\hat{H}_{S}$ and the reaction coordinate, is now coupled to a residual bath with the effect described by the new spectral function $J_{\rm RC}(\omega)=\sum_{k}|f_{k}|^{2}\delta(\omega-\omega_{k})$. The harmonic modes of the residual reservoir are described via the sets $\{\hat{b}_{k}\}$ with frequencies $\omega_{k}$, which are linear combinations of the original harmonic modes $\{\hat{c}_{k}\}$. Both $\lambda$ and $\Omega$ may be obtained from the original spectral function through the relations $\lambda^{2}=\frac{1}{\Omega}\int_{0}^{\infty}{\rm{d}}\omega\;\omega J(\omega)$ and $\Omega^{2}=\frac{\int_{0}^{\infty}{\rm{d}}\omega\;\omega^{3}J(\omega)}{\int_{0}^{\infty}{\rm{d}}\omega\;\omega J(\omega)}$ Iles-Smith et al. (2014).
@@
An effective Hamiltonian treatment may be employed in this model, which has been successful at describing equilibrium thermodynamics Anto-Sztrikacs et al. (2023a, b). We define a polaron transformation
@@
An analytical expression may be obtained for Eq. (7). We show in Appendix A that for the model in Eq. (2) we get
@@ II.3 Dynamics
We study the dynamics of the system by employing four tiers of approximations:
(i) The hierarchical equations of motion Tanimura (2020); Lambert et al. (2023), a numerically-exact method. Due to the increased complexity that arises as the system-bath interaction energy increases, with this method we are limited to interaction energies that are weak to intermediate. Details on the the HEOM are exposed in Appendix C.
@@
The QME for the dynamics of the matrix elements of the density matrix $\hat{\rho}_{\rm RC}(t)$ reads Nitzan (2013)
$\frac{{\rm d}\rho_{{\rm RC},mn}}{{\rm d}t}$ $=-\mathrm{i}\omega_{mn}\rho_{{\rm RC},mn}$
$+\sum_{jk}\left[R_{mj,jk}(\omega_{kj})\rho_{{\rm RC},kn}+R^{*}_{nk,kj}(\omega_{jk})\rho_{{\rm RC},mj}\right.$
$-\left.R_{kn,mj}(\omega_{jm})\rho_{{\rm RC},jk}+R^{*}_{jn,mk}(\omega_{km})\rho_{{\rm RC},jk}\right],$ (11)
@@
such that $\hat{H}^{\rm RC}_{S}\ket{\omega_{m}}=\omega_{m}\ket{\omega_{m}}$. In Eq. (II.3), the subindices denote the matrix elements in the energy eigenbasis while $R_{ab,cd}$ are the elements in this basis of the Redfield tensor. We have defined $\omega_{mn}=\omega_{m}-\omega_{n}$. The Redfield equation assumes that the dynamics begin from a system-bath factorized initial conditions. In arriving to Eq. (II.3), we note that the usual Born-Markov approximation has been carried out, but not the secular approximation. This means that, in general, Eq. (II.3) does not define a generator of a dynamical semigroup, and this may lead to non-positivity of the reduced density matrix along the time trajectory. This is typically not an issue at sufficiently weak coupling, which is guaranteed via our Markovian embedding in Sec. II.2. However, to confirm that the dynamics generated in our model under the conditions we have presented are correct, we have compared the time trajectories of observables computed from Eq. (II.3) with the numerically exact HEOM method. In Appendix C we demonstrate that when solving the Redfield QME on the RC Hamiltonian by following Eq. (II.3) we get the same results as with the hierarchical equations of motion.
@@
$R_{mn,jk}(\omega)$ $=\left[S_{\rm RC}\right]_{mn}\left[S_{\rm RC}\right]_{jk}\int_{0}^{\infty}{\rm d}\tau e^{\mathrm{i}\omega\tau}\langle\hat{B}(\tau)\hat{B}(0)\rangle,$ (13)
where $\hat{S}_{\rm RC}=\left(\hat{a}^{\dagger}+\hat{a}\right)$ is the system operator that couples to the bath in the enlarged Hamiltonian, $\hat{B}=\sum_{k}f_{k}(\hat{b}^{\dagger}_{k}+\hat{b}_{k})$ is the bath operator. The expectation value in Eq. (13) is taken with respect to the thermal state of the residual bath, i.e., $\langle\bullet\rangle=\mathrm{Tr}[\bullet\hat{\rho}_{B}]$, where $\hat{\rho}_{B}=e^{-\beta\hat{H}_{B}}/\mathrm{Tr}[e^{-\beta\hat{H}_{B}}]$, $\hat{H}_{B}=\sum_{k}\omega_{k}\hat{b}^{\dagger}_{k}\hat{b}_{k}$ and $\beta=1/T$ the inverse temperature. The integral can be readily evaluated to obtain
$R_{mn,jk}(\omega)$ $=\left[S_{\rm RC}\right]_{mn}\left[S_{\rm RC}\right]_{jk}\left[\Gamma_{\rm RC}(\omega)+\mathrm{i}\delta_{\rm RC}(\omega)\right],$ (14)
where $\Gamma_{\rm RC}(\omega)$ is the symmetric part of the correlation function and $\delta_{\rm RC}(\omega)$ is the so-called Lamb shift term, which we neglect in our calculations. We note that, in lack of a general solution for the diagonal form of $\hat{H}^{\rm RC}_{S}$, Eq. (II.3) is typically solved numerically with a truncated manifold of the reaction coordinate, such that only $M$ levels of the harmonic mode are kept to tractably study the asymptotic dynamics.
@@
The dynamics generated by the bath from the perspective of the effective Hamiltonian theory follows the same steps as above, with the only difference being that, as opposed to considering the eigenbasis of $\hat{H}^{\rm RC}_{S}$ for the Redfield tensor, one uses the eigenbasis of the effective Hamiltonian $\hat{H}^{\rm eff}_{S}$ from Eq. (9) along with the coupling operator $\hat{S}=\sum_{i=1}^{N}\hat{\sigma}^{x}_{i}$ for our particular choice in the spin model. Similarly, the simulation of the system dynamics at tier (iv), with the original Hamiltonian, proceeds based on Eq. (2).

### v2

@@ I Introduction
There exists, in turn, a present interest in understanding environmental effects in quantum systems. Such is the case in, for instance, biological systems and light-harvesting complexes Duan et al. (2022). Emerging technologies such as thermal machines Kosloff and Levy (2014); Goold et al. (2016); Benenti et al. (2017); Mitchison (2019); Josefsson et al. (2018); Ronzani et al. (2018); Mosso et al. (2019); Carrega et al. (2024), which operate in the quantum regime, also require a deep understanding of system-environment effects for efficient operation. Moreover, light-matter interactions between the vacuum fluctuations and a material placed inside the cavity provide a new scheme in controlling/engineering material/molecular properties Garcia-Vidal et al. (2021); Hübener et al. (2021); Owens et al. (2022); Schlawin et al. (2022); Bloch et al. (2022); Galego et al. (2015); Herrera and Spano (2016). An important remark is that in most of these cases, the system-environment interaction energy cannot be regarded as an asymptotically weak parameter in the configuration. This limits the applicability of standard perturbative open quantum systems theory Breuer and Petruccione (2007).
@@
Starting from out-of-equilibrium initial conditions, it is also of interest to understand the transient dynamics that lead the system and bath to the state of thermal equilibrium. In this respect, many impressive techniques have come to fruition in the last three decades to target the problem. These include; for example, path integral methods Makri (1995); Segal et al. (2010); Kilgour et al. (2019); Kundu and Makri (2023); Strathearn et al. (2018); Gribben et al. (2020), hierarchical equations of motion (HEOM) Tanimura (2020); Lambert et al. (2023), Monte Carlo approaches Erpenbeck et al. (2023); Härtle et al. (2015); Cohen et al. (2015), techniques based on machine learning calculations Luo et al. (2022), tensor networks Prior et al. (2010); Strathearn et al. (2018); Brenes et al. (2020); Cygorek et al. (2022), generalized quantum master equations Shi and Geva (2003); Zhang et al. (2006); Kelly et al. (2016), Davydov ansatze Chen and Zhao (2017); Zhao (2023); Zhou et al. (2016) and diverse Markovian embeddings Woods et al. (2014); Chin et al. (2010); Imamoglu (1994); Garraway (1997a, b); Menczel et al. (2024); all of which can be applied efficiently to study dynamics of system-bath configurations depending on the nature of the microscopic model, energy scales, and computational costs.
@@
The effective Hamiltonian theory Anto-Sztrikacs et al. (2023a, b) has been presented as an alternative to fully numerical techniques. Previous works, focused on the equilibrium and steady-state behavior of impurity systems coupled to bosonic thermal reservoirs, showed that this theory can provide an understanding of strong coupling effects in thermal and electronic devices without assumptions on the energy scale of the system-bath interaction energy. The theory relies on a Markovian embedding with a subsequent unitary transformation that imprints environmental effects onto the system’s degrees of freedom. One then considers a specific sub-manifold of the resulting Hamiltonian, and such an approximation can be justified under certain assumptions on the temperature and the bath’s spectral function. As a whole, the effective Hamiltonian mapping prepares a dressed system Hamiltonian that embeds strong coupling effects. The effective system itself is assumed to couple to a residual bath in a Born-Markov limit Breuer and Petruccione (2007). With this method, one can obtain analytical expressions for equilibrium states in the weak, intermediate, and strong system-bath coupling regimes Anto-Sztrikacs et al. (2023b). Furthermore, the technique can be employed even for complex quantum systems that may display phase transitions induced by collective or individual baths Min et al. (2023).
Here, we assess the capability of the effective Hamiltonian theory to capture and elucidate both the equilibrium and transient dynamics of system-bath configurations that are initially prepared out of equilibrium. Although this task can be addressed from e.g., numerical renormalization group (NRG) techniques Orth et al. (2010) and multi-D1 ansatze Deng et al. (2016) at zero temperature with semi-analytical arguments, here we focus on nonzero temperature baths.
@@
In Sec. II we describe our models, the theory of effective Hamiltonian, and the Redfield quantum master equation (QME) employed to describe the transient dynamics. Our results are divided into Sec. III and Sec. IV, devoted to equilibrium properties and dynamical signals, respectively. In light of our results, we provide a general assessment of the effective Hamiltonian method for both transient or steady-state problems in Sec. V. Concluding remarks and the experimental relevance of our results are included in Sec. VI.
II Method and bath-induced interactions
This Section is devoted to the description of the models, methods and techniques employed in our work. In Sec. II.1, we describe the open quantum system model we consider here. In Sec. II.2, we present the analytical effective Hamiltonian treatment used to study dynamics. The simulation method we use is explained in Sec. II.3. We also perform a spectral analysis of the effective Hamiltonian in Sec. II.4. This analysis is relevant for understanding the dynamics of out-of-equilibrium configurations.
@@ II.1 Spin-bath models
Figure 1: A two-spin system ($\hat{H}_{S}$) with spin splittings $\Delta_{1}$ and $\Delta_{2}$ is coupled to a thermal bath ($\hat{H}_{B}$) at inverse temperature $\beta$ with a coupling operator $\hat{S}$, modeled by an infinite collection of harmonic modes. The spectral density of the reservoir is given by $J(\omega)$. The spins are not coupled to each other but an effective coupling arises from the interaction with the common bath.
@@
The effect of the environment is typically described via a spectral density function, which may be defined as $J(\omega)=\sum_{k}|t_{k}|^{2}\delta(\omega-\nu_{k})$. The choice of $\hat{S}$ has intricate consequences for the dynamics of the system and, whenever the interaction energy between the system and environment is non-negligible, it will also affect the thermal equilibrium state Anto-Sztrikacs et al. (2023a). In this work, we shall focus on the case whereby the coupling to the environment generates coherent evolution towards the equilibration of the system, by choosing $\hat{S}$ along a different component than that of the spin-system Hamiltonian.
@@ II.2 Reaction-coordinate mapping and effective Hamiltonian theory
We want to make minimal assumptions about the system-bath interaction energy and, therefore, we cannot resort to the conventional weak-coupling open systems theory Breuer and Petruccione (2007). For certain spectral density functions of the bath, $J(\omega)$, it is beneficial to employ the reaction-coordinate mapping to go beyond the weak-coupling approximation Hughes et al. (2009a, b); Martinazzo et al. (2011); Iles-Smith et al. (2014, 2016); Strasberg et al. (2016); Hughes et al. (2009a); Anto-Sztrikacs and Segal (2021); Correa et al. (2019); Restrepo et al. (2019).
For the model at hand, we extract a single collective reaction coordinate (RC) from the bath to which the system is strongly coupled. This mapping is useful if the residual bath, i.e., the one composed of the degrees of freedom of the bath remaining after the extraction of the reaction coordinate, is weakly coupled to the enlarged system Hamiltonian. The enlarged system Hamiltonian is composed of the original system and the reaction coordinate. After such mapping, Eq. (1) translates to Hughes et al. (2009a, b); Martinazzo et al. (2011); Iles-Smith et al. (2014, 2016); Strasberg et al. (2016); Anto-Sztrikacs and Segal (2021); Anto-Sztrikacs et al. (2023a); Correa et al. (2019); Restrepo et al. (2019)
@@
In the above equation, the bosonic operators $\hat{a}$ and $\hat{a}^{\dagger}$ pertain to the reaction coordinate with frequency $\Omega$. The parameter $\lambda$ is the coupling between the original system $\hat{H}_{S}$ and the reaction coordinate via the system operator $\hat{S}$. The enlarged system, which is composed of $\hat{H}_{S}$ and the reaction coordinate, is now coupled to a residual bath with the effect described by the new spectral function $J_{\rm RC}(\omega)=\sum_{k}|f_{k}|^{2}\delta(\omega-\omega_{k})$. The harmonic modes of the residual reservoir are described via the sets $\{\hat{b}_{k}\}$ with frequencies $\omega_{k}$, which are linear combinations of the original harmonic modes $\{\hat{c}_{k}\}$. Both $\lambda$ and $\Omega$ may be obtained from the original spectral function through the relations $\lambda^{2}=\frac{1}{\Omega}\int_{0}^{\infty}{\rm{d}}\omega\;\omega J(\omega)$ and $\Omega^{2}=\frac{\int_{0}^{\infty}{\rm{d}}\omega\;\omega^{3}J(\omega)}{\int_{0}^{\infty}{\rm{d}}\omega\;\omega J(\omega)}$ Iles-Smith et al. (2014).
@@
An effective Hamiltonian treatment may be employed in this model, which has been successful at describing equilibrium thermodynamics Anto-Sztrikacs et al. (2023a, b). We define the so-called polaron transformation [see discussion below Eq. (10)]
@@
An analytical expression may be obtained for Eq. (7). We show in Appendix A that for the model in Eq. (2) the application of the polaron mapping and the truncation of the energy spectrum leads to
@@
The basic idea behind the polaron transformation is that the interaction between a quantum system (for example, electrons) and its environment (phonons) leads to the dressing of the quantum system by the environment, the former now referred to as a “polaron”. In the case of a charged particle, the coupling to the lattice modes effectively increases the mass of the electron, thus slowing it down.
The polaron transformation, which is unitary, is designed to (even if partially) decouple the quantum system from the environment. As an outcome, the system’s Hamiltonian becomes dressed by the bath. Studying quantum dynamics under a polaron picture is beneficial at strong system-bath coupling. This problem has been investigated in many works, with some examples in Refs. Nazir (2009); Lee et al. (2012); Nicolin and Segal (2011); Segal (2014). The advantage of the polaron picture is that in the so-called polaron frame, one can identify a perturbative parameter that replaces the original (nonpertuabtive) system-bath coupling  McCutcheon and Nazir (2010).
In our approach, the transformation in Eq. (4) shifts the reaction-coordinate Hamiltonian into the so-called polaron frame. As showed in Ref. Anto-Sztrikacs et al. (2023a), this step extends the range of applicability of the reaction-coordinate mapping.
@@ II.3 Dynamics
We study the dynamics of the system by employing four tiers of methods:
(i) The hierarchical equations of motion Tanimura (2020); Lambert et al. (2023), a numerically-exact method. While the method is exact, for practical purposes in numerical calculations one needs to truncate the hierarchy of equations and the number of terms pertaining to the expansion of the bath correlation function. Due to the increased complexity that arises as the system-bath interaction energy increases, with this method we obtain accurate results in the weak-to-intermediate interaction energy regime. Details on the the HEOM are presented in Appendix C.
@@
The QME for the dynamics of the matrix elements of the density matrix $\hat{\rho}(t)$ reads Nitzan (2013)
$\frac{{\rm d}\rho_{mn}}{{\rm d}t}$ $=-\mathrm{i}\omega_{mn}\rho_{mn}$
$-\sum_{jk}\left[R_{mj,jk}(\omega_{kj})\rho_{kn}+R^{*}_{nk,kj}(\omega_{jk})\rho_{mj}\right.$
$-\left.R_{kn,mj}(\omega_{jm})\rho_{jk}-R^{*}_{jm,nk}(\omega_{kn})\rho_{jk}\right],$ (11)
@@
such that $\hat{H}^{\rm RC}_{S}\ket{\omega_{m}}=\omega_{m}\ket{\omega_{m}}$. In Eq. (II.3), the subindices denote the matrix elements in the energy eigenbasis while $R_{ab,cd}$ are the elements in this basis of the Redfield tensor. We have defined $\omega_{mn}=\omega_{m}-\omega_{n}$. The Redfield equation assumes that the dynamics begin from a system-bath factorized initial conditions. In arriving to Eq. (II.3), we note that the usual Born-Markov approximation has been carried out, but not the secular approximation. To confirm that the dynamics generated in our model under the conditions we have presented are correct, specifically, the assumption of weak coupling and Markovianity to the residual bath, we have compared the dynamics of observables computed from Eq. (II.3) with the numerically exact HEOM method. In Appendix C we demonstrate that when solving the Redfield QME on the RC Hamiltonian by following Eq. (II.3) we get the same results as with the hierarchical equations of motion.
@@
$R_{mn,jk}(\omega)$ $=S_{mn}S_{jk}\int_{0}^{\infty}{\rm d}\tau e^{\mathrm{i}\omega\tau}\langle\hat{B}(\tau)\hat{B}(0)\rangle,$ (13)
where $\hat{S}=\left(\hat{a}^{\dagger}+\hat{a}\right)$ is the system operator that couples to the bath, with matrix elements $S_{mn}$ $[S_{jk}]$ in the energy eigenbasis of $\hat{H}^{\rm RC}_{S}$. $\hat{B}=\sum_{k}f_{k}(\hat{b}^{\dagger}_{k}+\hat{b}_{k})$ is the bath operator. The expectation value in Eq. (13) is taken with respect to the thermal state of the residual bath, i.e., $\langle\bullet\rangle=\mathrm{Tr}[\bullet\hat{\rho}_{B}]$, where $\hat{\rho}_{B}=e^{-\beta\hat{H}_{B}}/\mathrm{Tr}[e^{-\beta\hat{H}_{B}}]$, $\hat{H}_{B}=\sum_{k}\omega_{k}\hat{b}^{\dagger}_{k}\hat{b}_{k}$ and $\beta=1/T$ the inverse temperature. The integral can be readily evaluated to obtain
$R_{mn,jk}(\omega)$ $=S_{mn}S_{jk}\left[\Gamma_{\rm RC}(\omega)+\mathrm{i}\delta_{\rm RC}(\omega)\right],$ (14)
where $\Gamma_{\rm RC}(\omega)$ is the symmetric part of the correlation function and $\delta_{\rm RC}(\omega)$ is the so-called Lamb shift term, which we neglect in our calculations Correa and Glatthard (2023). We note that, in lack of a general solution for the diagonal form of $\hat{H}^{\rm RC}_{S}$, Eq. (II.3) is typically solved numerically with a truncated manifold of the reaction coordinate, such that only $M$ levels of the harmonic mode are kept to tractably study the dynamics.
@@
The dynamics generated by the bath from the perspective of the effective Hamiltonian theory follows the same steps as above, with the only difference being that, as opposed to considering the eigenbasis of $\hat{H}^{\rm RC}_{S}$ for the Redfield tensor, one uses the eigenbasis of the effective Hamiltonian $\hat{H}^{\rm eff}_{S}$ from Eq. (9) along with the coupling operator $\hat{S}=\sum_{i=1}^{N}\hat{\sigma}^{x}_{i}$ for our particular choice in the spin model.
Similarly, the simulation of the system dynamics at tier (iv), with the original Hamiltonian, proceeds based on Eq. (2) with the original spectral function $J(\omega)$.